Q1. What is the molarity of an ethanol solution, C2HsOH, when 65.5 grams of C2HsOH is dissolved in 675 mL of DI water? Q2. What mass, in grams, of potassium oxalate, K2C:20, is needed to prepare 625 mL of a 0.235 M K2C204 solution? Q3. A certain experiment requires 15.5 grams of sodium hydroxide, NaOH. Volume wise, how many milliliters (mL) of a 0.525 M solution of NaOH is needed?
Answer #1

1) molarity is defined as the number of moles of solute in 1 L of solvent.

moles of ethanol = mass/ molar mass

= 65.5g/ 46g/mol

= 1.424 mol

volume of solution = 675mL = 0.675L

Thus molarity of solution = 1.424/0.675

= 2.1095M

2) Molarity = 0.235M

molar mass of K2C2O4 = 166g/mol

mass of sample = x g

moles = xg/166g/mol

volume of solution = 625mL = 0.625L

Thus

0.235 M = (x/166) /0.625

thus mass of sample =24. 3815g

Q3)mass of NaOH = 15.5g

molarmass of NaoH = 40g/mol

volume of solution = ?

molarity = 0.525M

Thus 0.525mol/L = (15.5/40)mol /V

V, volume of solution = 0.738 L

= 738mL